OpenStudy (anonymous):

If X^6-Y=9999, what are the values of X and Y?

OpenStudy (anonymous):

The permissible values of X and Y that fit your constraint, will need to meet the following criterium: x^6 - y = 9999, so x^6 = 9999 + y, so x^6 -9999 = y. Every combination of (x,y) that fits this function is acceptable.